OpenStudy (insa):

6 stools and 4 chairs cost $58 but 5 stools and 2 chairs cost $35. find the cost of each stool and each chair.

OpenStudy (anonymous):

Your question is a case of linear equations in 2 variables. Let the no. of stools be x, and chairs be y. 6x + 4y = 58 5x + 2y = 35 Use the elimination method or substitution method, whichever you like. I will use elimination. Multiply equation 2 by 2 You now have: 6x + 4y = 58 10x + 4y = 70 SUBTRACT. You now have -4x + 0 = -12 4x = 12 x = 12/4 = 3 Substitute (x = 3) in the first equation. 6(3) + 4y = 58 18 + 4y = 58 4y = 58 - 18 = 40 y = 40/4 = 10 Therefore, each stool (x) costs $3 and each chair (y) costs $10. Hope this helps.

OpenStudy (insa):

thankyou so much....:)

OpenStudy (anonymous):

anytime